OpenStudy (anonymous):

simplify. 3^1/2 times 3^1/2

OpenStudy (anonymous):

\[3^{\frac{ 1 }{ 2 }} \times 3^{\frac{ 1 }{ 2}}\]

OpenStudy (anonymous):

Add the exponents, 1/2+1/2=1 so it is 3^1 or just 3

OpenStudy (anonymous):

thnx so much